OpenStudy (anonymous):

Find the reciprocal of the number. Multiply to check your answer. A. B. C. D.

OpenStudy (anonymous):

2/3 a.1/2 b.3/2 c.2 d.3

OpenStudy (anonymous):

please cn somone help me

OpenStudy (anonymous):

@Phebe

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@Megan30400

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

do u ind heling with somemore

OpenStudy (anonymous):

Find the reciprocal of the number. Multiply to check your answer. A. B. C. D.

OpenStudy (anonymous):

2 3/4 a.2 4/3 b.11/4 c.4/11 d.4/23

OpenStudy (anonymous):

i think it c

OpenStudy (anonymous):

or d

OpenStudy (anonymous):

c is correct

OpenStudy (anonymous):

reciprocal: an expression or function so related to another that their product is unity; the quantity obtained by dividing the number one by a given quantity. So you would do \[1 \div \frac{ 3}{ 4 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

3/5 divide 1/4

OpenStudy (anonymous):

how do u answer this then put it into fraction

OpenStudy (anonymous):

sorry do \[1 \div \frac{ 23}{ 4 } = \frac{ 4 }{ 23 }\]

OpenStudy (anonymous):

so the answer is 4/23

OpenStudy (anonymous):

yeah, to check do \[\frac{ 23}{ 4} \times \frac{ 4 }{ 23 } = 1\]

OpenStudy (anonymous):

wow thanx

OpenStudy (anonymous):

could i be 3/20

OpenStudy (anonymous):

most of the times with reciprocals just swap the numerator and the denominator

OpenStudy (anonymous):

bcs theres o opyions of 4/23

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

2 5/4

OpenStudy (anonymous):

oooooh

OpenStudy (anonymous):

I was doing the reciprocal question 2!

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

ok can i just give u a new qustion

OpenStudy (anonymous):

3/5 divide 1/4

OpenStudy (anonymous):

\[\frac{ 3 }{ 5} \div \frac{ 1}{ 4 } = \frac{ 3 }{ 5 } \times \frac{ 4 }{ 1} = \frac{ 12 }{ 5 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i understand thanx a lot

OpenStudy (anonymous):

it is okay, the best thing to do is have a go at some more examples :) practise works with maths :)

OpenStudy (anonymous):

ok will do thats sounds great